# Kilonova Light Curve Parameter Estimation Using Likelihood-Free Inference

### v2

@@ 3.2 Normalizing Flows
For our LFI, we leverage normalizing flows, which transform a simple initial density into a complex distribution by composing a series of learnable transformations (Jimenez Rezende & Mohamed, 2015; Durkan et al., 2020). We use a three-dimensional standard normal base distribution and Masked Affine Autoregressive Transforms to compose the normalizing flow, as this transformation only requires one forward pass calculation to evaluate the density (Papamakarios et al., 2017). The LFI is conditioned with the representation, $f(d)$, from Eq. (3.1) from the summary embedding space described in Section 3.1 and trained by maximizing the log-likelihood with an Adam optimizer. We halve the learning rate when the loss plateaus with a patience of 2 epochs to avoid stagnation.
@@
Only the shifted view is utilized to train the normalizing flow. 472,300 shifted light curves with unique noise and unique combinations of all 5 parameters from Table 1 comprise the training dataset. Hence, this method is suited to the realistic scenario where there is uncertainty in the arrival time, or the distance to the source is not known precisely. The representation, $f(d)$, ensures that this uncertainty is marginalized over. Our best performing model is composed of 9 blocks, 5 transforms, and 90 hidden layers, selected by hyperparameter tuning (see Section 3.3).
@@
Figure 4: Upper: Training (solid) and validation (dashed) loss curves for variations of the flexibility of the embedding network. The use of an untrained representation is shown in red. The dark blue, light blue, and green curves represent training the flow with a pretrained representation. In dark blue, we allow the representation network to train alongside the flow. In light blue, we allow the parameters in $f$ to train, while the $h$ layer remains fixed. In green, we freeze the weights of the representation network while training the flow. We find that keeping the representation network frozen results in the lowest and fastest convergence. Lower: Percentile-Percentile plots from 100 test light curves, with 1000 posterior samples from each drawn by the flow. The fraction of instances where the parameters lie within the confidence interval is plotted for the three inferred parameters, with $M_{ej}$in blue, $V_{ej}$in orange, and $X_{lan}$in green. Overlaid are the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals in decreasing opacity. We find that our samples lie within the confidence intervals, indicating that the correct true parameters are inferred consistently.
The loss curves displayed in Figure 4 demonstrate the negative log probability of several embedding states, all utilizing a fixed batch size of 25, learning rate of $0.0000921$, and the same training dataset. The learning rate, batch size, and configurable model parameters are determined through hyperparameter tuning. See Section 3.3 and Appendix A. The frozen case accepts a preconditioned embedding and freezes the weights, while the unfrozen case allows all parameters of the embedding to vary with the flow. For the partially frozen case, we only freeze the weights of the $h$ layer. We investigate the effect of tuning the embedding alongside the flow and find similar rates of convergence. This allows us to select a lighter model with fewer than 1 million configurable parameters. On the other hand, an untrained representation initialized with random weights converges slower.We conclude that the embedding is a valuable step in marginalizing the time and distance values and reducing the overall size of the model. For a model size of $\sim$ 800,000 parameters, training requires 50 minutes to train for 100 epochs on one NVidia V100 GPU.
@@
We reserve a small sample of 100 light curves for testing the flow and PyMultiNest. 1000 posteriors are drawn from the trained flow for each light curve, and we calculate fraction of events that lie within the credible interval. This is plotted in Figure 4, where we see that the test cases all lie within the credible $3\sigma$ bands. Our model consistently returns accurate inferences for all three parameters. Section 4 further discusses the resulting posteriors from PyMultiNest and the flow.
@@ 3.3 Hyperparameter Tuning
To achieve optimal configurations of network parameters, hyperparamenter tuning is implemented. We utilize the ray.tune Python package to determine batch sizes and learning rates for our training loops, as well as find ideal configurations of our model parameters (Liaw et al., 2018). We searched over 100 combinations of potential dimensionality, kernel sizes, number of $h$ layers, number of blocks, and number of final dimensions for our representation network $\Gamma$ for a maximum of 50 epochs, or steps. Similarly, we searched through 100 combinations of number of transforms, blocks, and hidden features for the normalizing flow over a maximum of 100 steps. The choices we used are displayed in Table 2, where options in square brackets are discrete and those in parenthesis are continuous. We use the Async HyperBand Scheduler from Ray to terminate unpromising trials in an effort to reduce the required GPU time, and the best run is selected based on the lowest average validation loss achieved. We utilize the informed tuning to finalize our selection of model parameters. See Appendix A for configuration specifics and loss metrics for the top ten hyperparameter trials.
@@
Learning Rate LogUniform($1\times 10^{-6}$, $1\times 10^{-4}$)
@@
Learning Rate LogUniform($1\times 10^{-6}$, $1\times 10^{-4}$)
@@ 4.1 Posteriors
Figure 5: Left panels: Fixed and shifted light curves for injection parameters of $\log_{10}(M_{ej}/M_{\odot})$$=-1.5$, $\log_{10}(V_{ej}/c)$$=-1.0$, $\log_{10}(X_{lan})$$=-6$. Upper left: Shifted light curve with $d_{L}=50$ Mpc. Bottom left: Shifted light curve with $d_{L}=200$ Mpc. Right panels: Corresponding posterior distributions with the shifted light curve posterior from the flow shown in red, its corresponding fixed counterpart inferred from the flow shown in blue, and a comparison to PyMultiNest shown in green. We see agreement between the fixed and shifted posteriors from the flow, and that the three inferred parameters are in agreement with the truth. For $M_{ej}$, we find similar constraints when comparing to PyMultiNest, but the flow further constrains $V_{ej}$and $X_{lan}$.
We compare the performance of our model to nested sampling via the package PyMultiNest (Buchner et al., 2014). We maintain the same priors used for data generation, shown in Table 1, while generating the posterior plots. As the normalizing flow intrinsically only knows the given data distribution, we expect its estimation to match the data generation priors. To similarly inform PyMultiNest of the distribution, we use the same prior, but allow the parameter search space to remain $10\%$ larger than the original distribution to accommodate values at the edges of our priors. In addition to PyMultiNest, we investigate the posteriors obtained by the sampler dynesty (Speagle, 2020) and conclude that the two methods produce equivalent posteriors. We utilize PyMultiNest for our comparison inferences as it operates faster.
We display our comparison of the fixed and shifted posteriors to PyMultiNest sampling in Figure 5 for one example injection of $\log_{10}(M_{ej}/M_{\odot})$, $\log_{10}(V_{ej}/c)$, and $\log_{10}(X_{lan})$. A comparison is also made between a signal at $d_{L}=50$ Mpc and one at $d_{L}=200$ Mpc, showing the difficulty nested sampling has when the light curve peak has shifted significantly. Additionally, distant light curves are dimmer and provide fewer data points for the sampling method to leverage. Our LFI method remains invariant time, as displayed by the complete overlapping of the fixed and the shifted posteriors. The distributions of the flow posteriors also remain consistent for both luminosity distances, demonstrating distance invariance.
To create a statistical sample of posteriors, we draw 100 combinations of our injection parameters from our uniform data distribution which serve as our test set. The light curve data is passed to PyMultiNest, which uses 2048 live points to sample a posterior result. We obtain posteriors from the flow utilizing the identical set of shifted light curves. The resulting estimates are compared for our three parameters of interest in Figure 6. We observe that the flow performs well for $\log_{10}(V_{ej}/c)$and $\log_{10}(M_{ej}/M_{\odot})$, while the inference for $\log_{10}(X_{lan})$is challenging. We expect to see a linear correlation between the true values and the inferences. In the case of $\log_{10}(X_{lan})$, the posterior is uninformative. We find that for all three parameters, the flow returns values within the true distribution found in the data, though the margin of error can lie outside of the true distribution. For PyMultiNest, $\log_{10}(V_{ej}/c)$remains a difficult parameter to constrain within the given data distribution. We notice similar behavior for $\log_{10}(X_{lan})$. In these cases, the nested sampler is unable to find the correct minima associated with the true parameter. We observe that the precision and accuracy of the two methods are therefore comparable for $\log_{10}(M_{ej}/M_{\odot})$. The flow has improved accuracy for $\log_{10}(V_{ej}/c)$, but cannot sufficiently provide reliable information for $\log_{10}(X_{lan})$.
@@
Figure 6: Plotted are 100 inferences and errors for $M_{ej}$, $V_{ej}$, and $X_{lan}$ from PyMultiNestin blue, and values from the normalizing flow in red. The inferred values of the parameter estimation methods are plotted against the true value of the parameter. We expect to see a linear pattern when estimation is accurate, as the inference will match the truth. Our model performs well for the quantities $\log_{10}(M_{ej})$ and $\log_{10}(V_{ej})$, but is uninformative for $\log_{10}(X_{lan})$. PyMultiNesthas occasional difficulties in constraining a value for $\log_{10}(V_{ej})$ and $\log_{10}(X_{lan})$ for the provided distributions. We utilize the identical shifted light curve sample used in the percentile-percentile plot in Figure 4 for these inferences.
@@ 4.2 Model Resilience
We investigate the results of interpolation to generate missing sections of data by removing four consecutive data points from each band near the peak and performing a linear interpolation over a 36 hour time span. The posteriors for the original light curve and the interpolated one are well aligned. A sample case is displayed in Figure 7 with the interpolated light curve shown on the left, and the posterior in purple overlaid with the posterior from the original light curve on the right. The two posteriors are well aligned, providing a good indication that linear interpolation can be sufficient when the data has missing values. Data can also be interpolated through the use of transformers to achieve uniform cadences from non-uniform detections in the case where linear interpolation is not adequate (Mishra-Sharma, 2024).
For further tests in lower cadence regimes, see Appendix B for details on using interpolation to enhance the rapidity of the cadence. We expect that observational light curve data will be observed at a non-uniform cadence, and it is possible that key points near the peak will go undetected. Our framework is capable of leveraging interpolation to produce a reliable parameter estimate.
Figure 7: Left: Original light curve data plotted in color with the interpolated points plotted as black stars for all three ZTF bands. Right: Posteriors for a light curve with $\log_{10}(M_{ej}/M_{\odot})$$=-1.1045$, $\log_{10}(V_{ej}/c)$$=-1.0095$, and $\log_{10}(X_{lan})$$=-6.2940$. Blue represents the flow results on the original data, and purple with four consecutive data points interpolated. The two posteriors overlap, indicating that interpolating these missing values has little effect on the resulting distribution.
@@ 5 Conclusion
We present a method to perform parameter estimation on kilonova lightcurves using LFI. Our framework provides a parameter estimation method competitive with pre-existing samplers. We demonstrate that self-supervised learning successfully marginalizes our invariant parameters of distance and time, decreases the dimensionality of our light curve data, and allows the normalizing flow to converge faster. We do note that our method can be improved in respect to estimating $X_{lan}$, which can be difficult to inference when distance is marginalized.
Missing data is a common problem that can limit many ML approaches to classification and inference. We demonstrate that for a full day of missing data, our model returns statistically similar inferences when linear interpolation is applied. Additionally, our model provides posteriors in under a second while PyMultiNest requires time on the order of minutes. This allows our model to be a viable alternative inference method that has the added benefit of time and distance invariance.
Our method is constrained by the choice of kilonova model. In general, light curve models are parametrized by different physical inputs. For example, the model described in Bulla (2023) allows for priors on wind, inclination angle, and electron fraction of the neutron stars in addition to the ejecta mass and velocity parameters found in the Ka2017 model. Including these physical parameters requires retraining and slight restructuring to our network. However, the approach remains consistent regardless of model choice.
@@
The use of self-supervised learning as a summarization technique and normalizing flows to perform LFI can also be applied to other light curve data. Previous work has investigated the role of machine learning techniques on supernova light curves. For example, the Superphot+ classifier uses a gradient-boosted machine architecture to assign types to ZTF supernova light curves (de Soto et al., 2024). Superphot+ relies on nested sampling to estimate several light curve parameters. With a novel dataset, our architecture could be applied to improve secondary classifiers by providing fast and accurate parameter estimation. Since we demonstrate flexibility in cadences and missing data in Section 4.2 and Appendix B, our framework is capable of being applied to observational data.

### v3

@@ 3.2 Normalizing Flows
For our LFI, we leverage normalizing flows, which transform a simple initial density into a complex distribution by composing a series of learnable transformations (Jimenez Rezende & Mohamed, 2015; Durkan et al., 2020). We use a three-dimensional standard normal base distribution and Masked Affine Autoregressive Transforms to construct the flow, as this transformation allows efficient, single-pass density evaluation (Papamakarios et al., 2017). The flow is conditioned on the representation, $f(d)$, from Eq. (3.1) which summarizes the input light curves as described in Section 3.1. The model is trained by minimizing the negative log-likelihood (i.e., $-\log\mathcal{L}(\phi)$ from Eq.( 2)) using the Adam optimizer. Additionally, we halve the learning rate when the loss plateaus with a patience of 2 epochs to avoid stagnation.
@@
Only the shifted view is utilized to train the normalizing flow. 472,300 shifted light curves with unique noise and unique combinations of the 5 parameters of Table 1 comprise the training dataset. Hence, this method is suitable for the realistic scenario where there is uncertainty in the arrival time or the distance to the source is not known precisely. The representation $f(d)$, ensures that this uncertainty is marginalized over. Our best performing model is composed of 9 blocks, 5 transforms, and 90 hidden layers, informed by a hyperparameter search (see Section 3.3).
@@
Figure 4: Upper: Training (solid) and validation (dashed) loss curves for variations of the flexibility of the embedding network. The use of an untrained representation is shown in blue. The pink, black, and green curves are pretrained embeddings with frozen, partially frozen ($f(d)$ flexible), and unfrozen parameters ($f(d)$ and $h(d)$ flexible), respectively. We find that allowing some flexibility of the representation network results in the fastest convergence and the most accurate parameter estimation. Lower: Percentile-percentile plots from 100 test light curves, each with 1000 posterior samples drawn from the flow. The fraction of instances where the parameters lie within the confidence interval is plotted for the three inferred parameters, with $\log_{10}(M_{\rm{ej}}/M_{\odot})$in blue, $\log_{10}(V_{\rm{ej}}/c)$in orange, and $\log_{10}(X_{\rm{lan}})$in green. Overlaid are the $1\sigma$, $2\sigma$, and $3\sigma$ confidence intervals in decreasing opacity. We find that our samples lie within the confidence intervals, indicating that the correct true parameters are inferred consistently.
The loss curves displayed in the upper panel of Fig. 4 demonstrate the negative log probability of several embedding states, all utilizing a fixed batch size of 25, learning rate of $9.2\times 10^{-5}$, and an identical training dataset. The learning rate, batch size, and configurable model parameters are determined through hyperparameter tuning. See Section 3.3 and Appendix A. To demonstrate the efficacy of the data summary, we track the effects of several variations of the summary informing our flow. In the frozen case, we precondition the embedding but freeze all trainable weights when it is relayed to the normalizing flow. The unfrozen case allows all parameters of the embedding to vary with the flow. For the partially frozen case, we only freeze the weights of the $h$ layer, leaving $f(d)$ flexible. From Fig. 3, we confirm that allowing partial flexibility enables further clustering of parameters and directly correlates to an increase in our estimation accuracy.
The upper panel of Fig. 4 further showcases the difference in convergence times of the various embedding states. We find that the partially frozen configuration converges the fastest, allowing us to select a lighter model with approximately 1 million configurable parameters. On the other hand, an untrained representation initialized with random weights converges slower. Though the untrained representation is capable of achieving lower losses, we find that the model returns uninformative posteriors and cannot correlate two light curves with similar intrinsic parameters. The trained embedding purposely maintains the structure of the latent space, preventing a further decrease in loss. We conclude that utlizing a data summary step is valuable in marginalizing the time and distance values and reducing the overall size of the model. For a model size of $\sim$ 1.2 million parameters, training requires 50 minutes for 100 epochs on one NVidia V100 GPU.
@@
We reserve a small sample of 100 light curves for testing the flow and PyMultiNest. 1000 posteriors are drawn from the trained flow for each light curve, and we calculate fraction of events that lie within the credible interval. This is plotted in the lower panel of Figure 4, where we see that the test cases all lie within the credible $3\sigma$ bands. Our model consistently returns accurate inferences for all three parameters. Additionally, diagonal lines indicate that the model has high confidence in the results it returns. Section 4 further discusses the resulting posteriors from PyMultiNest and the flow.
@@ 3.3 Hyperparameter Tuning
To achieve optimal configurations of network parameters, hyperparamenter tuning is implemented. We utilize the ray.tune Python package to determine batch sizes and learning rates for our training, as well as find ideal configurations of our model parameters (Liaw et al., 2018). We searched over 100 combinations of potential dimensionality, kernel sizes, number of $h$ layers, number of blocks, and number of final dimensions for our representation network $\Gamma$ for a maximum of 50 epochs, or steps. Similarly, we searched through 100 combinations of number of transforms, blocks, and hidden features for the normalizing flow over a maximum of 100 steps. The choices we used are displayed in Table 3, where options in square brackets are discrete and those in parenthesis are continuous. We use the Async HyperBand Scheduler from Ray to terminate unpromising trials in an effort to reduce the required GPU time, and the best run is selected based on the lowest average validation loss achieved. We utilize the informed tuning to finalize our selection of model parameters. See Appendix A for configuration specifics and loss metrics for the top ten hyperparameter trials. We also note that lower-dimensional models generally perform with a decreased accuracy and precision. In particular, we find that reducing the dimensionality of the embedding network disrupts the adjustments made by the normalizing flow, resulting in a preferred higher dimension of 7.
@@
Learning Rate LogUniform($10^{-6}$, $10^{-4}$)
@@
Learning Rate LogUniform($10^{-6}$, $10^{-4}$)
@@ 4.1 Posteriors
Figure 5: Left panels: Fixed and shifted light curves for injection parameters of $\log_{10}(M_{\rm{ej}}/M_{\odot})$$=-1.44$, $\log_{10}(V_{\rm{ej}}/c)$$=-0.81$, and $\log_{10}(X_{\rm{lan}})$$=-8.76$. Upper left: Shifted light curve with $d_{L}=50$ Mpc. Bottom left: Shifted light curve with $d_{L}=200$ Mpc. Right panels: Corresponding posterior distributions with the shifted light curve posterior from the flow shown in red, its corresponding fixed counterpart inferred from the flow shown in green, and a comparison to PyMultiNest shown in blue. We see agreement between the fixed and shifted posteriors from the flow, and that the three inferred parameters are in agreement with the truth. We find that the flow is consistent regardless of the distance, while PyMultiNest faces more difficulty constraining parameters at $d_{L}=200$ Mpc.
@@
Figure 6: Plotted are 77 inferences and errors for $M_{\rm{ej}}$, $V_{\rm{ej}}$, and $X_{\rm{lan}}$ from PyMultiNest in blue, and values from the normalizing flow in red. The inferred values of the parameter estimation methods are plotted against the true value of the parameter. The black line represents the behavior we expect where the inferred and true parameters are the same. Overall, our model performs well for the quantities $\log_{10}(M_{ej})$, $\log_{10}(V_{ej})$, and $\log_{10}(X_{lan})$. PyMultiNest has occasional difficulties in constraining a value for $\log_{10}(V_{ej})$ and $\log_{10}(X_{lan})$ for the provided distributions, while our method has difficulty for high $\log_{10}(M_{ej})$. We utilize the identical shifted light curve sample used in the percentile-percentile plot in the lower panel of Figure 4 for these inferences.
We compare the performance of our model to nested sampling via the package PyMultiNest (Buchner et al., 2014). We maintain the same priors used for data generation, shown in Table 1, while generating the posterior plots. To inform PyMultiNest of the distribution, we use the same prior, but allow the parameter search space to remain $10\%$ larger than the original distribution to accommodate values at the edges of our priors. We assume the same priors for our model. The default sampler used by NMMA for light curve analysis is PyMultiNest, thus we utilize this sampler for our comparison inferences.
We display our comparison of the fixed and shifted posteriors to PyMultiNest sampling in Figure 5 for one example injection of $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$. A comparison is made between a signal at $d_{L}=50$ Mpc and one at $d_{L}=200$ Mpc, showing the difficulty nested sampling has when the light curve peak has shifted significantly. Additionally, distant light curves are dimmer and provide fewer data points for the sampling method to leverage. Our LFI method remains invariant to time, as displayed by the complete overlapping of the fixed and the shifted posteriors. The distributions of the flow posteriors also remain consistent for both luminosity distances, demonstrating distance invariance.
To create a statistical sample of posteriors, we draw 100 combinations of our injection parameters from our uniform data distribution which serve as our test set. The light curve data is passed to PyMultiNest, which uses 2048 live points to sample a posterior result. Due to an error in PyMultiNest’s sampling software that returns the same initial value for $\log_{10}(V_{\rm{ej}}/c)$, 23 out of the 100 test cases are discarded from the sample. The error in PyMultiNest’s software is outside of the scope of this paper. We obtain posteriors from the flow utilizing the identical set of 77 shifted light curves. The resulting estimates are compared for our three parameters of interest in Figure 6.
We expect to see a linear correlation between the true values and the inferences. We observe that the flow performs well for all three parameters: $\log_{10}(M_{\rm{ej}}/M_{\odot})$, $\log_{10}(V_{\rm{ej}}/c)$, and $\log_{10}(X_{\rm{lan}})$. Certain areas of parameter space are notably challenging to accurately predict. In particular, small quantities of lanthanide can be misjudged by both the normalizing flow and by PyMultiNest. High amounts of lanthanides have a distinct effect in the reddening of a kilonova light curve. As that amount reduces, the effect on the observed light curve is minimal. As a result, smaller values of $\log_{10}(X_{lan})$ are increasingly difficult to predict. We note that our recovery of $\log_{10}(V_{ej}/c)$ has an increased precision than PyMultiNest, while the spread of $\log_{10}(M_{ej}/M_{\odot})$ is similar between the two methods.
Table 4: Accuracy of the two methods. We check whether the true value falls within the predicted result’s range, defined by its error bars for both the flow and PyMultiNest. We then calculate the percentage of the 77 results that contain the true value within the error.
Parameter PyMultiNest LFI
$\log_{10}(M_{\rm{ej}}/M_{\odot})$ 62.34 % 66.23 %
$\log_{10}(V_{\rm{ej}}/c)$ 58.44 % 64.94 %
$\log_{10}(X_{\rm{lan}})$ 61.04 % 72.73 %
We compute the percentage of accurate inferences from both methods in Table 4. A result is deemed accurate if the truth falls within the associated error bars surrounding the estimate. Combining this information with the precision distributions in Figure 6, we conclude that our method is competitive with PyMultiNest in both accuracy and precision for all three parameters of interest.
@@ 4.2 Model Resilience
As we intend for our framework to be used with a variety of optical instruments, it is imperative that the model can adjust to several cadences. For our first test, we discard every other detection and linearly interpolate new data points in their stead. We find that the model has sufficient information to generate an inference with high accuracy. Figure 7 displays a sample posterior of the interpolation in grey with the original in red alongside the respective interpolated light curve. We observe alignment between the two posteriors. This test indicates that our model has some flexibility to accommodate different observational cadences to perform accurate parameter estimation.
We expect observational light-curve data to be observed at a nonuniform cadence, and it is possible that key points near the peak will go undetected. Data can also be interpolated through the use of transformers to achieve uniform cadences from nonuniform detections in the case where linear interpolation is not adequate (Mishra-Sharma, 2024). Similar work on analyzing supernova light curves has utilized Gaussian Process fitting to interpolate unevenly sampled light curves as a data pre-processing step prior to using normalizing flows (Villar, 2022). These methods provide avenues for further investigation into the creation of a more robust model for telescopic data.
For further tests in lower cadence regimes, see Appendix B for details on using interpolation to replace missing portions of the data.
Figure 7: Left: Original light curve data plotted in color with the interpolated points overplotted as black stars for all three ZTF bands. Right: Posteriors for a light curve with $\log_{10}(M_{\rm{ej}}/M_{\odot})$$=-1.0417$, $\log_{10}(V_{\rm{ej}}/c)$$=-1.3917$, and $\log_{10}(X_{\rm{lan}})$$=-7.0289$. Blue represents the flow results on the original data, and yellow with every other data point interpolated. The two posteriors overlap, indicating that interpolating these missing values has little effect on the resulting distribution.
@@ 5 Conclusion
We present a method to perform parameter estimation on kilonova lightcurves using LFI. Our framework provides a parameter estimation method competitive with pre-existing samplers. We demonstrate that self-supervised learning successfully marginalizes our invariant parameters of distance and time, decreases the dimensionality of our light curve data, and allows the normalizing flow to converge faster. Additionally, our model provides posteriors in less than a second, while PyMultiNest requires time on the order of minutes. This allows our model to be a viable alternative inference method that has the added benefit of time and distance invariance.
Missing data is a common problem that can limit many ML approaches to classification and inference. We demonstrate that for a full day of missing data, our model returns statistically similar inferences when linear interpolation is applied. Further investigation into utilizing observationally realistic light curves will be required to launch this model for analysis on data from Rubin.
Our method is constrained by the choice of kilonova model. In general, light curve models are parametrized by different physical inputs. For example, the model described in Bulla (2023) allows for priors on wind, inclination angle, and the electron fraction of the neutron star in addition to the ejecta mass and velocity parameters found in the Ka2017 model. Including these physical parameters requires retraining and slight restructuring to our network. However, the approach remains consistent regardless of model choice.
@@
The use of self-supervised learning as a marginalization technique and normalizing flows to perform LFI can also be applied to other light curve parameters. Previous work has investigated the role of machine learning techniques on supernova light curves. For example, the Superphot+ classifier uses a gradient-boosted machine architecture to assign types to ZTF supernova light curves (de Soto et al., 2024). Superphot+ relies on nested sampling to estimate several light curve parameters. With a novel dataset, our architecture could be applied to improve secondary classifiers by providing fast and accurate parameter estimation. Since we demonstrate flexibility in cadences and missing data in Section 4.2 and Appendix B, our framework is capable of being applied to observational data.